ISIS Logbooks General
  General, Page 1 of 9  Not logged in ELOG logo
Entry  Fri Apr 15 15:06:09 2016, Thomas Planche, , , I1/ISIS tunning 

15:03: scanning duty cycle and recording I1 voltages/currents, data saved under: xstrip$ddir:[SAVE]TP_20160415_I1tunning.

15:08: taking an EE scan with 6A arc, EInzel lens=28kV, PE=3.40 V, for comparison with last year's data. Data saved under the same file.

Entry  Wed Apr 13 17:12:47 2016, Thomas Planche, , , Optics box skimmer bias test 

Arc current set to 7 A, PE=3.40 V, gas flow=2.7cc,

Took a first EE scan for refrefence.

 

TW#74 (Neg. PS, cable#YCB2-TOP) connected to C6.

TW#72 (Neg. PS, cable#4002) connected to C5.


TW#73 (Pos. PS, cable#4000) connected to C9.

 

17:20: scanning TW#72. A few -100 V on #72 reduced Einzel lens current from 250 uA down to 40 uA.

17:23: set #72 to -400V, taking EE scan

17:27: #72 back to 0V, scanning #73. Does no good. Reduces beam on BeamStop, increased PulserDump.

17:31: #73 set to +400 V, taking EE scan.

17:38: #73 set to +2kV, #54 went down ~250 uA, could get ~200 uA back by changing steerer #43 (vertical) by 100 V.

17:43: #73 back to 0. Voltage readback is not 0, but 22 V.

17:44: now scanning #74. At -30V Einzel lens current increased by 50 ua (no change after that), no change on BeamStop.

 

Quick conclusion: Negative bias on C5 reduces a lot Einzel lens current. Pos. bias on C9 only make things worse.  Neg. bias on C6 has almost no effect.

 

 

 

Entry  Tue Apr 28 12:51:36 2015, Thomas Planche, , , Optics box steerer #44 not properly connected 

Robert L. said the HV cable SHV connector was not inserted all the way through the I1:44 feedthrough.

Entry  Tue Apr 28 11:25:09 2015, Raymond Dube, Routine, General, Added 1 Meg. ballast resistor to I1:45 HV PS TIS0550D-rev15.pdf

The Pulser moves the beam around and something loads the I1:45 HV PS use to set the Bias voltage for Collimators C5, C6, & C9.

With a setting of 170 volts the PS output would vary from 170 volts to 270 volts, to prevent I1:45 PS output from moving about we installed a 1 Meg/12.5W resistor, so you will draw 100uA/100 volts of Bias.

I updated TIS0550D rev.15 to reflect the changes to I1:45.

 

Entry  Tue Apr 14 13:25:19 2015, Thomas Planche, , , I1 electron filter magnet rough measurment 

Very quickly done measurement of the magnetic field along the beam axis through the I1 electron filter magnet:

z [mm] B [G]
0 -120
10 -210
20 +60
30 +200
40 60

Note1: z=0 corresponds to the tip of Keerthi probe aligned with the upstream edge of the extraction electrode.

Note2: orientation of the probe was kept ~constant. Initial orientation was found by maximazing the readback at z=40 mm.

Note3: error bar is huge. Measuremenet done very quickly. Something like +/- 20%.

Entry  Fri Mar 13 14:39:20 2015, John Chow, Routine, General, IB Bender 3 & 4 HV relays IB-B3B4-HV-Relay_2015.jpg

To day we inspected the IB B3 & B4 HV Relays, W/permit # C2015-03-13-7,  the small copper wires between the HV insulator feedthroughs and the HV relay contacts was replaced on all 4 HV relays with a 30KV rated wire, see photo: IB-B3B4-HV-Relay_2015.jpg

 

 

Entry  Fri Jan 16 09:39:43 2015, Thomas Planche, , , I1 study with Cyc. OFF and fixed PE voltage 

Data taken with steerer #43 set at 200 V:

Note: double peak observed for arc current higher than 4 A. Einzel (60) = 28.0 kV

1 2 3 4 5 6 7 8 9 10 11 12 13
 Arc current [A]  Arc volatage  PE volatge [V]  EE voltage [kV]  EE current [mA]  Bias current [uA]  Einzel current [uA]  H2 valve [cc]  BST 54 [uA]

PE cur.

[A]

Coll.[51]

[uA]

 Bias [45]

[uA]

 Beam dump [53] [uA]

 1.0  110  3.40  2.029  2.37  340  10  1.90  107  0.95  40  3.9  7
1.5  110  3.40  1.551  2.43  600  21  2.00  256  1.34  100  15  7
 2.0  110  3.40  1.375  2.61  770  40  2.10  377  1.71  150  51  12
 3.0  110  3.40  1.316  3.40  1120  64  2.30  488  2.42  310  86  34
 4.0  110  3.40  1.341  4.54  1450  85  2.5  531  3.08  480  118  53
 4.0  110  3.40  1.521  4.98  1500  70  2.5  533  3.08  520  96  63
 5.0  110  3.40  1.394  5.45  1750  105  2.75  537  3.69  680  155  57
 5.0  110  3.40  1.658  6.48  1890  84  2.75  594  3.69  630  111  77
 6.0  110  3.40  1.460  6.48  2030  114  3.00  527  4.30  820  176  99
 6.0  110  3.40  1.761  7.30  2180  92  3.00  629  4.30  760  126  80
 7.0  110  3.40  1.514  7.33  2280  110  3.20  518  4.88  1000  182  81
 7.0  110  3.40  1.846  8.26  2480  103  3.20  654  4.88  910  143  94
 8.0  110  3.40  1.578  8.29  2510  114  3.40  516  5.45  1000  192  115
 8.0  110  3.40  1.924  9.32  2760  116  3.40  660  5.45  1000  157  108
 9.0  110  3.40  1.629  9.14  2720  111  3.60  512  5.98  1000  188  110
 9.0  110  3.40  1.993  9.85  3020  120  3.60  662  5.98  1000  171  140

 

Entry  Wed Jan 14 10:23:00 2015, Thomas Planche, , , I1 study with cyc. OFF -- second day 

Data taken with steerer #43 OFF (vertical steerer before Einzel lens):

 Arc current [A]  Arc volatage  PE volatge [V]  EE voltage [kV]  EE current [mA]  Bias current [uA]  Einzel current [uA]  H2 valve [cc]  BST 54 [uA]
 2.0  110  3.44  1.316  3.37  830  50  2.10  395
 3.0  110  2.89  1.372  7.71  1260  70  2.40  451
 4.0  110  2.93  1.548  10.52  1640  72  2.50  484
 5.0  110  3.44  1.690  7.65  1950  84  2.70  522
 6.0  110  3.71  1.788  6.48  2210  100  2.90  550
 7.0  110  3.83  1.861  5.42  2430  113  3.20  563
 8.0  110  3.99  1.922  5.68  2650  117  3.30  568

Data taken with steerer #43 set at 200 V:

 Arc current [A]  Arc volatage  PE volatge [V]  EE voltage [kV]  EE current [mA]  Bias current [uA]  Einzel current [uA]  H2 valve [cc]  BST 54 [uA]
 2.0  110  3.56  1.343  3.40  830  47  2.10  416
 3.0  110  3.40  1.306  4.48  1190  72  2.40  502
 4.0  110  3.36  1.363  6.53  1540  91  2.50  545
 5.0  110  3.95  1.678  5.51  1840  82  2.70  596
 6.0  110  3.41  1.768  6.59  2210  97  2.90  633
 7.0  110  3.79  1.851  5.51  2440  107  3.20  654
 8.0  110  3.40  1.951  9.08  2860  116  3.50  666
 9.0  110  3.24  1.866  13.57  3020  122  3.50  727

Data taken with steerer #43 set at 200 V. These are 'wrong' data but I do not want to though away data...

 5.0  110  3.05  1.433  10.46  1890  84  2.70  563
 6.0  110  3.01  1.495  12.31  2170  130  2.90  565
Entry  Tue Jan 13 13:18:34 2015, Thomas Planche, , , I1 study with cyclotron OFF xt51ps.pngxt51ps.pngxt51ps.pngxt51ps.pngxt51ps.png

The table below gives a summary; for more detailes, look at the attached files.

 Arc current [A]  Arc volatage  PE volatge [V]  EE voltage [kV]  EE current [mA]  Bias current [uA]  Einzel current [uA]  H2 valve [cc]  BST 54 [uA]
 2.0  110  3.01  1.350  4.42  790  40  2.00  365
 4.0  110  3.01  1.426  12.31  1.51  79  2.20  457
 6.0  110  4.02  1.729  4.84  2.0  92  2.80  533
 4.0  110  2.93  1.473  9.90  1.55  78  2.50  473
 8.0  110  4.18  1.868  5.1  2.46  114  3.20  561
Entry  Mon Jan 12 15:01:48 2015, Raymond Dube, Routine, General, Test with Negative PS for Einzel Lens 

Last December 19th 2014 we had a development shift to run the Einzel lens with a Negative power supply  (-15KV 20mA) unfortunately we could not run the Emittance scanner but the conclusion was that we were better off with a Positive Einzel lens power supply, it causes less aberration.  

Entry  Mon Oct 6 16:06:08 2014, Raymond Dube, Routine, General, Source 1 ISO-TX info ISO-TX-I1.pdfTIS0550D-rev14.pdf

I added the Isolation Transformer to the Source 1 Electrical Diagram, ref: TIS0550D rev. 14 and I scanned a copy of Jaroslav order for the Isolation Transformer "ISO-TX-I1.pdf"

Entry  Tue Sep 23 16:36:53 2014, Raymond Dube, Routine, General, VIB Skimmers connections TIS0517D-rev3.pdf

I updated the drawing "TIS0517D rev.3" for the VIB Skimmers connections. 

Note:  The IB & VIB Skimmers Bias Voltage comes from a Glassman Power Supply (+1KV 500uA) in CON.007 not from the Skimmer Bias Controller in CON.009 - the PS output is set today at + 200 Volts

there is a voltage read back (TW# IB:579)  from the SKimmer Bias Controller (10V = 10,000 V) but it is not very accurate, it reads 225 volts.  We should read the Bias voltage directly from the Glassman P/S.

Entry  Tue Sep 23 16:26:50 2014, Raymond Dube, Routine, General, Coll. & BST batteries voltage inspection Bias-Box-Collimators_Voltage.pdf

This morning I checked the voltage on the 300V batteries for the Collimators and Beamstops, see attached file: Bias-Box-Collimators_Voltage.pdf

Note 1: There are some devices that use 2 batteries or 600 volts: COL 007, COL 017, BST 030, BST 003, BST 013 & BST 023

Note 2:  We need to pay attention to voltage on the battery for COL 061 because it draws beam current and drains the battery

Note 3:  These 300 Volts batteries are obsolete

Entry  Tue Jun 3 14:17:52 2014, Thomas Planche, , , Possition offset on BPM 12 and 16 

Alex Sorokin swapped connectors of BPM 12 and 16. I recorded the beam position shown on page 5BPM before and after the swap:

   before swap[mm] after[mm]  average [mm]
BPM 12 X  -0.70  2.51  0.91
BPM 12 Y  2.51  1.33  1.92
BPM 16 X  -0.32  -0.25  -0.29
BPM 16 Y  0.48  0.06  0.54

 

 

Entry  Mon May 12 12:00:53 2014, Raymond Dube, Routine, General, New Optics Box Devices Col1-YCB1-Pulser.jpgI1-XCB-Old_New.jpgI1-YCB-Old_New.jpgPulser_Dump-2014.JPG

Attached are some photos of new devices installed in the Optics Box during the last shutdown.

Entry  Tue May 6 19:48:12 2014, Suresh, , , Emittance scan 

 

WS WS scan time Bias volt [#COL-45] Emit-Vertical [D.F = 10 %]
61 18:44:45 600 SES_140506_183832.dat
61 18:42:56 0 SES_140506_185113.dat
030 18:46:15 600 SES_140506_183832.dat
030 18:46:47 0 SES_140506_185113.dat

 

Entry  Tue May 6 13:41:50 2014, Raymond Dube, Routine, General, New SS flexible line for H2 flow I1-H2-BOP.JPG

Today Tim replaced the soft copper line inside the BOP for the Hydrogen bottle with a flexible SS line to make the H2 bottle replacement easier and also to relieve stress on the Quick connector and prevent H2 leakage.

Entry  Tue May 6 13:35:51 2014, Raymond Dube, Routine, General, Updates to TIS0806D & TIS0550D TIS0550D-rev13.pdfTIS0806D-rev6.pdf

I updated drawings:  TIS0860D rev.6 and TIS0550D rev.13 to reflect the changes with the Negative power supplies for I1:72 and I1:74.

Entry  Mon May 5 16:00:17 2014, Raymond Dube, Routine, General, Development Shift, I1:72 & 74 Neg. PS 2014-05-05-12_24_31.WS030_rondbeam-2.png

During Development shift today we replaced the Positive PS (+3KV 5mA) for I1:72 & I1:74 with Negative PS ( - 3KV 5mA) this allowed us the reduce by half the steering on I1:73 and I1:76 both PS are Positive 3KV 5mA. 

The effect of these changes was a more round beam at WSM:030 which was the desired effect.

Entry  Tue Apr 29 08:33:25 2014, Raymond Dube, Routine, General, Source 1 down time 28//04/2014 

Yesterday the Extraction lens current and voltage were not stable, we had to stop beam and cancel the Proton therapy, there was 2 patients left.  We found a damaged Hydrogen line (H2 bottle was empty, 0 PSI) we replaced the H2 bottle and replaced the "Back Plate, new filament" and the Source ran fine after that.

PS:  We should be thinking about a 2 bottle set up for the Hydrogen supply.

ELOG V2.9.2-2455